OpenStudy (anonymous):

What is the exact value of squared root of 32 over 5 square root of 14

OpenStudy (anonymous):

\[\sqrt{\frac{32}{5} }\times \sqrt{14}\] is this what you mean?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its\[\sqrt{32}/ 5\sqrt{14}\]

OpenStudy (anonymous):

ah ok thats 4.233

OpenStudy (anonymous):

or do you mean \[\frac{8\sqrt{7}}{5}\]

OpenStudy (anonymous):

yes, i figures where i was messing up,

OpenStudy (redneckoutlaw):

i have 3 more

OpenStudy (redneckoutlaw):

What is the exact value of 14 times z squared all over the square root of z to the ninth power.?

OpenStudy (redneckoutlaw):

@makayla_finkbeiner_